OpenStudy (anonymous):

which equation show x^2+6x-2=0 rewritten by completing the square?

OpenStudy (unklerhaukus):

(x+a)^2 + b = (x+a)(x+a) + b = x^2 + 2ax + a^2 + b compare the second term ( 2ax ) with the second term ( 6x ) in the expression x^2 + 6x - 2 what is a?